Gaussian distribution, bell shaped distribution, normal distribution, all are same. The values needed for the given distribution of temperatures are:
- The fraction of patients expected to have a fever greater than 103.58°F is 0.0026(probability) or 0.26%(percent) or 26/10000(fraction) approx
- The fraction of patients expected to have a temperature between 101.43°F and 102.74°F is 0.7487 (probability) or 74.87%(percent) or 7487/10000(fraction) approx.
How to get the z scores?
If we've got a normal distribution, then we can convert it to standard normal distribution and its values will give us the z score.
If we have
\(X \sim N(\mu, \sigma)\)
(X is following normal distribution with mean \(\mu\)standard deviation \(\sigma\))
then it can be converted to standard normal distribution as
\(Z = \dfrac{X - \mu}{\sigma}, \\\\Z \sim N(0,1)\)
(Know the fact that in continuous distribution, probability of a single point is 0, so we can write
\(P(Z \leq z) = P(Z < z) )\)
Also, know that if we look for Z = z in z tables, the p value we get is
\(P(Z \leq z) = \rm p \: value\)
For the given case, let the random variable X be tracking the temperature readings(in degree Fahrenheit) of flu patients for given clinic in given condition. Then, by the specified data, we get:
\(X \sim N(102.00, 0.566)\)
The fraction of patients can be expressed in percentage or probability.
Calculating the needed fractions:
a) Fraction of patients having temperature > 103.58°F = P(X > 103.58°F)
Using the standard normal distribution transformation, we get:
P(X > 103.58°F) = \(P(Z > \dfrac{103.58 - \mu}{\sigma}) = P(Z > \dfrac{103.58 - 102}{0.566}) \approx P(Z > 2.79)\)
P(Z > 2.79) = 1 - P(Z ≤ 2.79)
From z tables, for z = 2.79, the p value is
Thus,
P(X > 103.58°F) = 1 - P(Z ≤ 2.79) = 1 - 0.9974= 0.0026
b) The fraction of patients expected to have a temperature between 101.43°F and 102.74°F = P(101.43°F≤ X ≤ 102.74°F)
P(101.43°F≤ X ≤ 102.74°F) = P( X ≤ 102.74°F) - P(X < 101.43°F)
Using standard normal distribution and z tables, we get:
\(P(X \leq 102.74) = P(Z \leq \dfrac{102.74 - 102}{0.566}) \approx P(Z < 1.31) =0.9049\)
and
\(P(X < 101.43) = P(X \leq 101.43) = P(Z \leq \dfrac{101.43 -102}{0.566}) \approx P(Z \leq -1.01)\\\\P(X < 101.43) \approx 0.1562\)
Thus,
P(101.43°F≤ X ≤ 102.74°F) = P( X ≤ 102.74°F) - P(X < 101.43°F)
= 0.9049 - 0.1562 = 0.7487
Thus,
The values needed for the given distribution of temperatures are:
- The fraction of patients expected to have a fever greater than 103.58°F is 0.0026(probability) or 0.26%(percent) or 26/10000(fraction) approx
- The fraction of patients expected to have a temperature between 101.43°F and 102.74°F is 0.7487 (probability) or 74.87%(percent) or 7487/10000(fraction) approx.
Learn more about standard normal distribution here:
https://brainly.com/question/10984889